# False signatures of non-ergodic behavior in disordered quantum many-body systems

### v2

@@ IV Artifacts in disorder averaged observables
Figure 7: The disorder averaged rescaled energy expectation value for the Néel state, $\epsilon_{\psi}$, and the associated standard deviation with disorder realization (shown here by the error bars ) varying with disorder strength for different system sizes. For the HSC Hamiltonian with RD (a) and QP disorder (b). Both (a) and (b) consider $2500$ disorder realizations.
@@
Figure 8: The probability distribution of site-resolved magnetization of eigenstates, $P(s_{l}^{z})$, for different system sizes and energies around $\epsilon=0.3$, the approximate average energy of the Néel state for both the RD (a) and QP (b) case. Both (a) and (b) consider $100$ eigenstates for at least $250$ disorder realizations with $W=0.5$. The bottom row provides a comparison with the distributions for site-resolved magnetization at long times obtained from dynamics of the initial Néel state, $P(\overline{s}_{l}^{z})$ again for both RD (c) and QP (d) disorder cases. Both (c) and (d) consider $2500$ disorder realizations.
@@ V Effects on time-dynamics
In this section, we examine the experimental consequences of the anomalous distributions of $s_{l}^{z}$ in eigenstates. We consider a system with Hamiltonian $H$, initialized in state |$\ket{\phi}$ and denote the time-evolved state by |$\ket{\phi(t)}=e^{-iHt}\ket{\phi}$. Long time average $\overline{\mathcal{O}}$ of expectation value of an observable $\hat{O}$ is given as
@@
where $c_{n}=\braket{n}{\phi}$. Performing the time integral, we note that only the diagonal terms, $m=n$, yield a non-vanishing contribution, showing that $\overline{\mathcal{O}}$ is given by the expectation value in the diagonal ensemble,
$\overline{\mathcal{O}}=\sum_{n}\braket{n}{\hat{\mathcal{O}}}{n}|c_{n}|^{2},$ (17)
defined by the initial state |$\ket{\phi}$. When the $H$ is ergodic and local, and the initial state is a product state, the energy fluctuations are subsextensive [4], $\delta E_{\phi}/E_{\phi}\sim 1/\sqrt{L}$, where $E_{\phi}=\braket{\phi}{H}{\phi}$ and $\delta E_{\phi}=\sqrt{\braket{\phi}{\hat{H}^{2}}{\phi}-E_{\phi}^{2}}$. In that case, the diagonal ensemble expectation value, (17), may be close to the microcanonical average $\mathcal{O}(E_{\phi})$, (1), at energy $E_{\phi}$. In an ergodic disordered system, the mechanism leading to the false signatures of non-ergodicity in the distributions of expectation values of local observables in eigenstates may lead to a non-trivial spread of the long-time averages, Eq. (16), of the local observables. In the following, we analyze this effect on the example of disordered HSC, showing how to assess the ergodicity of the system correctly.
We fix the initial quench state as the Néel state $|\ket{\psi}\equiv\ket{\uparrow\downarrow\cdots\uparrow\downarrow}$ and follow the time evolution for individual disorder realizations using Chebyshev propagation scheme [66] - see Appendix B. For disorder strength $W=0.5$, the average rescaled energy of |$\ket{\psi}$ in the HSC model with both the RD and QP disorder is $\braket{\psi}{\hat{H}_{\mathrm{HSC}}}{\psi}\equiv$ $\epsilon_{\psi}\approx 0.3$, c.f. Fig. 7. The latter value is significantly distant from the rescaled energy $\braket{\epsilon}\approx 0.62$ corresponding to $\braket{E}=-1/4$. At the considered system size, the expectation value $\hat{s}_{l}^{z}$ enters a stationary regime already at time-scales $O(100)$. Hence, as an approximation for the infinite time average in Eq. (16) for the operator $\hat{s}^{z}_{l}=\hat{\mathcal{O}}$, we consider its long time average as
@@
Figure 9: The probability distribution of long time average of the site-resolved magnetization, $\overline{s}_{l}^{z}$, (18), corrected by the microcanonical expectation value, $s_{l}^{z}(E_{{\color[rgb]{0.5,0.5,0}{{\psi}}}})$, in the Néel state, $\mathcal{P}(\overline{s}_{l}^{z}-s_{l}^{z}(E_{{\color[rgb]{0.5,0.5,0}{{\psi}}}}))$, for the HSC Hamiltonian with RD for different system sizes. The microcanonical expectation is calculated as an average over $10$ eigenstates, and the probability distributions are found by considering $2500$ disorder realizations.
The distributions of $\overline{s}^{z}_{l}$, obtained from the simulated time dynamics for different sites $l$ and disorder realizations, do not follow a naive expectation of their narrow concentration around the mean value for disordered ergodic systems. As shown in Fig. 9, the microcanonical ensemble expectation value, $s^{z}_{l}(E_{\psi})$, concentrates around the the long-time average value for the Néel state, $\overline{s}^{z}_{l}$, up to corrections decreasing with system size $L$. Similarly to the case of eigenstates considered earlier, the significant spread of the long-time averages, $\overline{s}^{z}_{l}$, may lead to false conclusions about breakdown of ergodicity. A remedy for this artifact may be, however, readily constructed, even at the single disorder realization level. Instead of analyzing the long-time dynamics for individual spin, $s^{z}_{l}(t)=\braket{\psi}{\hat{s}_{l}^{z}(t)}{\psi}$, one may subtract from it the diagonal ensemble expectation given by (16). Without access to eigenstates, this expectation may be simply taken as a long-time average, defined above, $\overline{s}^{z}_{l}$. The result of such a procedure can be seen in Fig. 10. The top row presents different $s^{z}_{l}(t)$ for a single disorder realization, showing the residual, spread of the saturation values, $\overline{s}^{z}_{l}$, present even at long times. When $\tilde{s}^{z}_{l}(t)=s^{z}_{l}(t)-\overline{s}^{z}_{l}$ are plotted, the spread dramatically narrows down, and $\tilde{s}^{z}_{l}(t)$ fluctuates, at long times, around zero mean, as one could expect for the ergodic dynamics.
Figure 10: The time evolution of the site-resolved magnetizations with the Néel state $s_{l}^{z}(t)=\braket{\psi}{\hat{s}_{l}^{z}(t)}{\psi}$ for the HSC Hamiltonian with RD, (a) and QP (b) disorders for a single disorder realization with $W=0.5$ and the system size $L=20$. The bottom row shows the same dynamics when the long time average of each spin, $\overline{s}_{l}^{z}$, is subtracted, $\tilde{s}_{l}^{z}\equiv s_{l}^{z}(t)-\overline{s}_{l}^{z}$.
@@ VI Conclusion
In this study, we have highlighted the subtle issues related to self-averaging in disorder-averaged observables, focusing on the experimentally relevant probe of site-resolved magnetization in spin-$1/2$ chains. Taking as an example the paradigmatic Heisenberg spin-$1/2$ chain, we systematically investigated various types of onsite disorder in the ergodic regime. Our results show that the nontrivial structures observed in the probability distributions of matrix elements of local observables arise as a direct consequence of the energy dependence of the microcanonical expectation value, which varies for each specific disorder realization. Importantly, we demonstrated that the expected Gaussian behavior predicted by random matrix theory can be recovered either by selecting eigenstates near the infinite-temperature energy window or by explicitly removing the mean microcanonical contribution. This work clarifies how misleading signatures of non-ergodic behavior can emerge and provides practical guidance for correctly interpreting numerical and experimental results in disordered quantum systems.
@@ Acknowledgements.
A.S.A., P.S., and J.Z. would like to acknowledge insightful discussions with Titas Chanda in the initial stages of this work. The work of A.S.A. and J.Z. was partially funded by the National Science Centre, Poland, project 2021/03/Y/ST2/00186 within the QuantERA II Programme (DYNAMITE) that has received funding from the European Union Horizon 2020 research and innovation programme under Grant agreement No 101017733 and by the National Science Centre, Poland, project 2021/43/I/ST3/01142 – OPUS call within the WEAVE programme. A.E.K. acknowledges the support of the National Science Centre, Poland, via Project No. 2021/42/A/ST2/00017. P.S. acknowledges fellowship within the “Generación D” initiative, Red.es, Ministerio para la Transformación Digital y de la Función Pública, for talent attraction (C005/24-ED CV1), funded by the European Union NextGenerationEU funds, through PRTR. L.V. acknowledges support from the Slovenian Research and Innovation Agency (ARIS), Research core funding Grants No. P1-0044, No. N1-0273, and No. J1-50005. We gratefully acknowledge the Polish high-performance computing infrastructure PLGrid (HPC Centers: ACK Cyfronet AGH) for providing computer facilities and support within computational grant no. PLG/2025/018400.
@@
Let us start by decomposing the HSC Hamiltonian with onsite fields as a sum of $2-$site and $1-$site local operators. Leading to the expression for the Hamiltonian as $\hat{H}_{\mathrm{HSC}}\equiv\sum_{i=1}^{L-1}[\hat{J}_{i,i+1}+\hat{V}_{i,i+1}]+\sum_{i=1}^{L}\hat{h}_{i}$, where $\hat{J}_{i,i+1}\equiv\frac{1}{2}(\hat{s}_{i}^{+}\hat{s}_{i+1}^{-}+\hat{s}_{i}^{-}\hat{s}_{i+1}^{+})$, $\hat{V}_{i,i+1}=\hat{s}_{i}^{z}\hat{s}_{i+1}^{z}$ and $\hat{h}_{i}\equiv h_{i}\hat{s}_{i}^{z}$. Considering the zero magnetization Fock basis in the $Z-$basis, which has a dimension $D=L!/((L/2)!(L/2)!)$, we will estimate the various traces of interest associated with the derivations of $\gamma_{0}$ and $\gamma_{1}$. That is, $\mathrm{Tr}[\hat{s}_{l}^{z}]$, $\mathrm{Tr}[\hat{H}_{\mathrm{HSC}}]$, $\mathrm{Tr}[\hat{H}_{\mathrm{HSC}}^{2}]$ and $\mathrm{Tr}[\hat{s}_{l}^{z}\hat{H}_{\mathrm{HSC}}]$.
@@
Working with the $Z-$base, it is immediately apparent that $\mathrm{Tr}[\hat{s}_{l}^{z}]$ is simply zero, so we start by calculating the trace of the Hamiltonian, $\mathrm{Tr}[\hat{H}_{\mathrm{HSC}}]$. The tunneling terms, $\hat{J}_{i,i+1}$, here will only have off-diagonal elements and thus do not contribute to the Hamiltonian trace. From $\mathrm{Tr}[\hat{s}_{l}^{z}]=0$ it is evident that the contributions from the onsite fields $\hat{h}_{i}$ are also zero. And the only terms that contribute are the interaction terms $\hat{V}_{i,i+1}$ which from the combinatorial restrictions associated with choosing two spins oriented in up/down directions within this magnetization restricted basis make a contribution to the basis as $\mathrm{Tr}[\hat{J}_{i,i+1}]=-\mathcal{D}/(4(L-1))$. Thus, considering all terms for the HSC Hamiltonian with open boundary conditions, one gets: $\mathrm{Tr}[\hat{H}_{\mathrm{HSC}}]=-D/4$.
Next, considering the higher momentum of the Hamiltonian, $\mathrm{Tr}[\hat{H}_{\mathrm{HSC}}^{2}]$, one needs to consider pure squares of the local operators of the Hamiltonian and products of different local operators. The terms of the tunneling part that lead to products with other local operators are once again purely off-diagonal and do not contribute, so one is left with terms of the form $\hat{J}_{i,i+1}^{2}$ or $\hat{J}_{i,i+1}\hat{J}_{j,j+1}$ (with $i\neq j$) , here again all terms of the second form will only have off-diagonal elements and will have zero contribution. On expanding the trace associated with $\hat{J}_{i,i+1}^{2}$ one finds $\hat{J}_{i,i+1}^{2}=\mathrm{Tr}[(\hat{s}_{i}^{x})^{2}(\hat{s}_{i+1}^{x})^{2}+(\hat{s}_{i}^{y})^{2}(\hat{s}_{i+1}^{y})^{2}+(\hat{s}_{i}^{x})(\hat{s}_{i+1}^{x})(\hat{s}_{i}^{y})(\hat{s}_{i+1}^{y})+(\hat{s}_{i}^{y})(\hat{s}_{i+1}^{y})(\hat{s}_{i}^{x})(\hat{s}_{i+1}^{x})]$. The first two terms here are just scalars and will just have the contribution $\mathrm{Tr}[(\hat{s}_{i}^{x})^{2}(\hat{s}_{i+1}^{x})^{2}+(\hat{s}_{i}^{y})^{2}(\hat{s}_{i+1}^{y})^{2}]=D/8$, while for the second term with the Pauli commutation relations one can recover a local operator of the form $\hat{V}_{i,i+1}$; then making use of the combinatorial restrictions one finds $\mathrm{Tr}[(\hat{s}_{i}^{y})(\hat{s}_{i+1}^{y})(\hat{s}_{i}^{x})(\hat{s}_{i+1}^{x})]=D/(8(L-1))$. On summing over all terms one finds the contribution from the tunneling part as $\mathrm{Tr}[\sum_{i,j}\hat{J}_{i,i+1}\hat{J}_{j,j+1}]=DL/8$. While considering the interaction part the cross-term product with the onsite field operators will have no contribution as it will just produce $3-$site or $1-$site $\hat{s}_{l}^{z}$ operators, both of which have vanishing trace. Now, considering the pure products of the interaction part, one has terms of the form $\mathrm{Tr}[\hat{V}_{i,i+1}\hat{V}_{j,j+1}]$. For $i=j$ the trace will give the scalar contribution $D/16$. While, for $|i-j|=1$ the trace will take a form $\frac{1}{4}\mathrm{Tr}[\hat{s}_{m}^{z}\hat{s}_{n}^{z}]$ that makes a contribution $-D/(16(L-1))$. And for $|i-j|>1$ one has the full $4-$site operator of the form $\mathrm{Tr}[\hat{s}_{i}^{z}\hat{s}_{i+1}^{z}\hat{s}_{j}^{z}\hat{s}_{j+1}^{z}]$ that again, making use of combinatorics associated with evaluating the probability of choosing different choices of spins within this constrained basis, gives $\mathrm{Tr}[\hat{s}_{i}^{z}\hat{s}_{i+1}^{z}\hat{s}_{j}^{z}\hat{s}_{j+1}^{z}]=\frac{3{D}}{16(L-1)(L-3)}$. Summing over all indices the total contribution arising from the interacting part is $\mathrm{Tr}[\sum_{i,j}\hat{V}_{i,i+1}\hat{V}_{j,j+1}]=L/8-\frac{1}{16(L-1)}[2(L-2)]+\frac{3{D}}{16(L-1)(L-3)}[(L-2)(L-3)]$. And finally considering the field terms we find traces of the form $\mathrm{Tr}[h_{i}\hat{s}_{i}^{z}h_{j}\hat{s}_{j}^{z}]$ which when $i=j$ is a trace over a scalar that gives $Dh_{i}^{2}/4$ while for $i\neq j$ gives terms similar to $\hat{V}$ and giving the trace value, $\frac{Dh_{i}h_{j}}{4(L-1)}$. On summing over indices, $\mathrm{Tr}[\sum_{i,j}\hat{h}_{i}\hat{h}_{j}]=+1/4\sum_{i=1}^{L}h_{i}^{2}-1/(4(L-1))\sum_{i\neq j}h_{i}h_{j}$. Collating all contributions from the three terms of the Hamiltonian one finds: $\mathrm{Tr}[\hat{H}_{\mathrm{HSC}}^{2}]/D=(3L^{2}-3L-1)/(16(L-1))+1/4\sum_{i=1}^{L}h_{i}^{2}-1/(4(L-1))\sum_{i\neq j}h_{i}h_{j}$.
@@
Finally, we look at $\mathrm{Tr}[\hat{H}\hat{s}_{j}^{z}]$. This trace just has a trace over local operators, which is up to a constant already dealt with while evaluating $\mathrm{Tr}[\hat{H}_{\mathrm{HSC}}^{2}]$. Here, there are no contributions from the tunneling and the interaction part based on insights mentioned above. While from the field operators, one gets a trace over operators of the form $\mathrm{Tr}[h_{i}\hat{s}_{i}^{z}\hat{s}_{j}^{z}]$, which have already been evaluated up to a scalar above; one finds in the case $i=j$ a contribution of $Dh_{i}/4$, while for $i\neq j$ the contribution is $-(\mathcal{D}h_{i})/(4(L-1))$. Summing over the indices, we find: $\mathrm{Tr}[\hat{H}_{\mathrm{HSC}}\hat{s}_{j}^{z}]/\mathcal{D}=1/4h_{j}-1/(4(L-1))\sum_{i\neq j}h_{i}$.
@@
Figure 11: The minimum and maximum energy for disorder averaged HSC Hamiltonian for both the RD and QP disorder (a,b) with different system sizes, $L$, for $W=0.5$. While the rescale energy, $\epsilon$, is shown at the average energy point, $\braket{\epsilon}$, corresponding to the energy point $E=0$, and the energy point of maximal DoS with varying $L$ for $W=0.5$ for the HSC Hamiltonian with RD (c) and QP (d) disorder. All curves here consider at least $4500$ disorder realizations.

### v3

@@ IV Artifacts in disorder averaged observables
Figure 7: The disorder averaged rescaled energy expectation value for the Néel state, $\epsilon_{\psi}$, and the associated standard deviation with disorder realization (shown here by the error bars ) varying with disorder strength for different system sizes. For the HSC Hamiltonian with RD (a) and QP disorder (b). Both (a) and (b) consider 250025002500 disorder realizations.
@@
Figure 8: The probability distribution of site-resolved magnetization of eigenstates, $P(s_{l}^{z})$, for different system sizes and energies around $\epsilon=0.3$, the approximate average energy of the Néel state for both the RD (a) and QP (b) case. Both (a) and (b) consider 100100100 eigenstates for at least 250250250 disorder realizations with $W=0.5$. The bottom row provides a comparison with the distributions for site-resolved magnetization at long times obtained from dynamics of the initial Néel state, $P(\overline{s}_{l}^{z})$ again for both RD (c) and QP (d) disorder cases. Both (c) and (d) consider 250025002500 disorder realizations.
@@ V Effects on time-dynamics
In this section, we examine the experimental consequences of the anomalous distributions of $s_{l}^{z}$ in eigenstates. We consider a system with Hamiltonian $\hat{H}$, initialized in state |$\ket{\phi}$ and denote the time-evolved state by |$\ket{\phi(t)}=e^{-i\hat{H}t}\ket{\phi}$. Long time average $\overline{\mathcal{O}}$ of expectation value of an observable $\hat{O}$ is given as
@@
where $c_{n}=\braket{n|\phi}$. Performing the time integral, we note that only the diagonal terms, $m=n$, yield a non-vanishing contribution, showing that $\overline{\mathcal{O}}$ is given by the expectation value in the diagonal ensemble,
$\overline{\mathcal{O}}=\sum_{n}\braket{n|\hat{\mathcal{O}}|n}|c_{n}|^{2},$ (17)
defined by the initial state |$\ket{\phi}$. When the $\hat{H}$ is ergodic and local, and the initial state is a product state, the energy fluctuations are subsextensive [4], $\delta E_{\phi}/E_{\phi}\sim 1/\sqrt{L}$, where $E_{\phi}=\braket{\phi|\hat{H}|\phi}$ and $\delta E_{\phi}=\sqrt{\braket{\phi|\hat{H}^{2}|\phi}-E_{\phi}^{2}}$. In that case, the diagonal ensemble expectation value, (17), may be close to the microcanonical average $\mathcal{O}(E_{\phi})$, (1), at energy $E_{\phi}$. In an ergodic disordered system, the mechanism leading to the false signatures of non-ergodicity in the distributions of expectation values of local observables in eigenstates may lead to a non-trivial spread of the long-time averages, Eq. (16), of the local observables. In the following, we analyze this effect on the example of disordered HSC, showing how to assess the ergodicity of the system correctly.
We fix the initial quench state as the Néel state $|\ket{\psi}\equiv\ket{\uparrow\downarrow\cdots\uparrow\downarrow}$ and follow the time evolution for individual disorder realizations using Chebyshev propagation scheme [66] - see Appendix B. For disorder strength $W=0.5$, the average rescaled energy of |$\ket{\psi}$ in the HSC model with both the RD and QP disorder is $\braket{\psi|\hat{H}_{\mathrm{HSC}}|\psi}\equiv$ $\epsilon_{\psi}\approx 0.3$, c.f. Fig. 7. The latter value is significantly distant from the rescaled energy $\braket{\epsilon}\approx 0.62$ corresponding to $\braket{E}=-1/4$. At the considered system size, the expectation value $\hat{s}_{l}^{z}$ enters a stationary regime already at time-scales $O(100)$. Hence, as an approximation for the infinite time average in Eq. (16) for the operator $\hat{s}^{z}_{l}=\hat{\mathcal{O}}$, we consider its long time average as
@@
Figure 9: The probability distribution of long time average of the site-resolved magnetization, $\overline{s}_{l}^{z}$, (18), corrected by the microcanonical expectation value, $s_{l}^{z}(E_{{\psi}})$, in the Néel state, $\mathcal{P}(\overline{s}_{l}^{z}-s_{l}^{z}(E_{{\psi}}))$, for the HSC Hamiltonian with RD for different system sizes. The microcanonical expectation is calculated as an average over 101010 eigenstates, and the probability distributions are found by considering 250025002500 disorder realizations.
The distributions of $\overline{s}^{z}_{l}$, obtained from the simulated time dynamics for different sites $l$ and disorder realizations, do not follow a naive expectation of their narrow concentration around the mean value equal to zero for disordered ergodic systems. As shown in Fig. 9, these distributions are centered around the microcanonical ensemble expectation value, $s^{z}_{l}(E_{\psi})$ calculated at the energy $E_{\psi}$ of the initial Néel state, up to corrections decreasing with system size $L$. This microcanonical ensemble expectation value is estimated as an average of $s^{z}_{l}$ over $N$ eigenstates with eigenenergies close to a desired energy $s^{z}_{l}(E)=1/N\sum_{n}\braket{n|\hat{s}^{z}_{l}|n}$.
Similarly to the case of eigenstates considered earlier, the significant spread of the long-time averages, $\overline{s}^{z}_{l}$, may lead to false conclusions about the breakdown of ergodicity. A remedy for this artifact may be, however, readily constructed, even at the single disorder realization level. Instead of analyzing the long-time dynamics for individual spin, $s^{z}_{l}(t)=\braket{\psi|\hat{s}_{l}^{z}(t)|\psi}$, one may subtract from it the microcanonical expectation. The result of such a procedure can be seen in Fig. 10. The top row presents different $s^{z}_{l}(t)$ for a single disorder realization, showing the residual spread of the saturation values, $\overline{s}^{z}_{l}$, present even at long times. When $\tilde{s}^{z}_{l}(t)=s^{z}_{l}(t)-s^{z}_{l}(E_{\psi})$ are plotted, the spread dramatically narrows down, and $\tilde{s}^{z}_{l}(t)$ fluctuates, at long times, around zero mean, as one could expect for the ergodic dynamics.
Figure 10: The time evolution of the site-resolved magnetizations with the Néel state $s_{l}^{z}(t)=\braket{\psi|\hat{s}_{l}^{z}(t)|\psi}$ for the HSC Hamiltonian with RD, (a) and QP (b) disorders for a single disorder realization with $W=0.5$ and the system size $L=20$. The bottom row shows the same dynamics when the microcanonical expectation for each spin, ${s}_{l}^{z}(E_{\psi})$ (obtained using 100 eigenstates) is subtracted, $\tilde{s}_{l}^{z}\equiv s_{l}^{z}(t)-{s}_{l}^{z}(E_{\psi})$.
@@ VI Conclusion
In this study, we have highlighted the subtle issues related to self-averaging in disorder-averaged observables, focusing on the experimentally relevant probe of site-resolved magnetization in spin-1/21/21 / 2 chains. Taking as an example the paradigmatic Heisenberg spin-1/21/21 / 2 chain, we systematically investigated various types of onsite disorder in the ergodic regime. Our results show that the nontrivial structures observed in the probability distributions of matrix elements of local observables arise as a direct consequence of the energy dependence of the microcanonical expectation value, which varies for each specific disorder realization. Importantly, we demonstrated that the expected Gaussian behavior predicted by random matrix theory can be recovered either by selecting eigenstates near the infinite-temperature energy window or by explicitly removing the mean microcanonical contribution. This work clarifies how misleading signatures of non-ergodic behavior can emerge and provides practical guidance for correctly interpreting numerical and experimental results in disordered quantum systems.
@@ Acknowledgements.
A.S.A., P.S., and J.Z. would like to acknowledge insightful discussions with Titas Chanda in the initial stages of this work. We thank also Konrad Pawlik for a careful reading of the manuscript. The work of A.S.A. and J.Z. was partially funded by the National Science Centre, Poland, project 2021/03/Y/ST2/00186 within the QuantERA II Programme (DYNAMITE) that has received funding from the European Union Horizon 2020 research and innovation programme under Grant agreement No 101017733 and by the National Science Centre, Poland, project 2021/43/I/ST3/01142 – OPUS call within the WEAVE programme. A.E.K. acknowledges the support of the National Science Centre, Poland, via Project No. 2021/42/A/ST2/00017. P.S. acknowledges fellowship within the “Generación D” initiative, Red.es, Ministerio para la Transformación Digital y de la Función Pública, for talent attraction (C005/24-ED CV1), funded by the European Union NextGenerationEU funds, through PRTR. L.V. acknowledges support from the Slovenian Research and Innovation Agency (ARIS), Research core funding Grants No. P1-0044, No. N1-0273, and No. J1-50005. We gratefully acknowledge the Polish high-performance computing infrastructure PLGrid (HPC Centers: ACK Cyfronet AGH) for providing computer facilities and support within computational grant no. PLG/2025/018400.
@@
Let us start by decomposing the HSC Hamiltonian with onsite fields as a sum of 22-2 -site and 11-1 -site local operators. Leading to the expression for the Hamiltonian as $\hat{H}_{\mathrm{HSC}}\equiv\sum_{i=1}^{L-1}[\hat{J}_{i,i+1}+\hat{V}_{i,i+1}]+\sum_{i=1}^{L}\hat{h}_{i}$, where $\hat{J}_{i,i+1}\equiv\frac{1}{2}(\hat{s}_{i}^{+}\hat{s}_{i+1}^{-}+\hat{s}_{i}^{-}\hat{s}_{i+1}^{+})$, $\hat{V}_{i,i+1}=\hat{s}_{i}^{z}\hat{s}_{i+1}^{z}$ and $\hat{h}_{i}\equiv h_{i}\hat{s}_{i}^{z}$. Considering the zero magnetization Fock basis in the $Z-$basis, which has a dimension $D=L!/((L/2)!(L/2)!)$, we will estimate the various traces of interest associated with the derivations of $\gamma_{0}$ and $\gamma_{1}$. That is, $\mathrm{Tr}[\hat{s}_{l}^{z}]$, $\mathrm{Tr}[\hat{H}_{\mathrm{HSC}}]$, $\mathrm{Tr}[\hat{H}_{\mathrm{HSC}}^{2}]$ and $\mathrm{Tr}[\hat{s}_{l}^{z}\hat{H}_{\mathrm{HSC}}]$.
@@
Working with the $Z-$base, it is immediately apparent that $\mathrm{Tr}[\hat{s}_{l}^{z}]$ is simply zero, so we start by calculating the trace of the Hamiltonian, $\mathrm{Tr}[\hat{H}_{\mathrm{HSC}}]$. The tunneling terms, $\hat{J}_{i,i+1}$, here will only have off-diagonal elements and thus do not contribute to the Hamiltonian trace. From $\mathrm{Tr}[\hat{s}_{l}^{z}]=0$ it is evident that the contributions from the onsite fields $\hat{h}_{i}$ are also zero. And the only terms that contribute are the interaction terms $\hat{V}_{i,i+1}$ which from the combinatorial restrictions associated with choosing two spins oriented in up/down directions within this magnetization restricted basis make a contribution to the basis as $\mathrm{Tr}[\hat{J}_{i,i+1}]=-{D}/(4(L-1))$. Thus, considering all terms for the HSC Hamiltonian with open boundary conditions, one gets: $\mathrm{Tr}[\hat{H}_{\mathrm{HSC}}]=-D/4$.
[…]
@@
Finally, we look at $\mathrm{Tr}[\hat{H}\hat{s}_{j}^{z}]$. This trace just has a trace over local operators, which is up to a constant already dealt with while evaluating $\mathrm{Tr}[\hat{H}_{\mathrm{HSC}}^{2}]$. Here, there are no contributions from the tunneling and the interaction part based on insights mentioned above. While from the field operators, one gets a trace over operators of the form $\mathrm{Tr}[h_{i}\hat{s}_{i}^{z}\hat{s}_{j}^{z}]$, which have already been evaluated up to a scalar above; one finds in the case $i=j$ a contribution of $Dh_{i}/4$, while for $i\neq j$ the contribution is $-({D}h_{i})/(4(L-1))$. Summing over the indices, we find: $\mathrm{Tr}[\hat{H}_{\mathrm{HSC}}\hat{s}_{j}^{z}]/{D}=1/4h_{j}-1/(4(L-1))\sum_{i\neq j}h_{i}$.
@@
Figure 11: The minimum and maximum energy for disorder averaged HSC Hamiltonian for both the RD and QP disorder (a,b) with different system sizes, $L$, for $W=0.5$. While the rescale energy, $\epsilon$, is shown at the average energy point, $\braket{\epsilon}$, corresponding to the energy point $E=0$, and the energy point of maximal DoS with varying $L$ for $W=0.5$ for the HSC Hamiltonian with RD (c) and QP (d) disorder. All curves here consider at least 450045004500 disorder realizations.